Give exact and approximate solutions to three decimal places for x^2 + 6x + 9 = 1
Answer by richwmiller(17219)

x^2 + 6x + 9 = 1
x^2 + 6x + 8=0
(x+2)(x+4)=0
x=-2
x=-4